# T2AcmrTempora-TLF

### v2

@@ 6 Conclusion and Future Work
The evaluation results show that this approach substantially improved the model’s factual knowledge, and the understanding of low-resource languages. It also shows that after such fine-tuning, the model’s responses are much better in terms of correctness and soundness, as assessed by native speakers and LLM as a judge.
In future work, we plan to apply this methodology to other languages and dialects. We further aim to work towards streamlining and automating the process as much as possible. We will also focus more on the modeling part of the experiments, and open-source culturally and institutionally relevant models for low-resource languages including Kazakh.
@@ 8 Ethics
  • Alyafeai et al. (2024) Zaid Alyafeai, Khalid Almubarak, Ahmed Ashraf, Deema Alnuhait, Saied Alshahrani, Gubran Abdulrahman, Gamil Ahmed, Qais Gawah, Zead Saleh, Mustafa Ghaleb, Yousef Ali, and Maged Al-shaibani. 2024. CIDAR: Culturally relevant instruction dataset for arabic. In Findings of the Association for Computational Linguistics: ACL 2024, pages 12878–12901, Bangkok, Thailand. Association for Computational Linguistics.
  • Anthropic (2024) Anthropic. 2024. Claude ai: An overview. https://www.anthropic.com.
  • Bandarkar et al. (2023) Lucas Bandarkar, Davis Liang, Benjamin Muller, Mikel Artetxe, Satya Narayan Shukla, Donald Husa, Naman Goyal, Abhinandan Krishnan, Luke Zettlemoyer, and Madian Khabsa. 2023. The belebele benchmark: a parallel reading comprehension dataset in 122 language variants.
@@
  • Brief et al. (2024) Meni Brief, Oded Ovadia, Gil Shenderovitz, Noga Ben Yoash, Rachel Lemberg, and Eitam Sheetrit. 2024. Mixing it up: The cocktail effect of multi-task fine-tuning on llm performance – a case study in finance.
@@
  • Cahyawijaya et al. (2024) Samuel Cahyawijaya, Holy Lovenia, Fajri Koto, Rifki Putri, Wawan Cenggoro, Jhonson Lee, Salsabil Akbar, Emmanuel Dave, Nuurshadieq Nuurshadieq, Muhammad Mahendra, Rr Putri, Bryan Wilie, Genta Winata, Alham Aji, Ayu Purwarianti, and Pascale Fung. 2024. Cendol: Open instruction-tuned generative large language models for Indonesian languages. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 14899–14914, Bangkok, Thailand. Association for Computational Linguistics.
@@
  • Cahyawijaya et al. (2023) Samuel Cahyawijaya, Holy Lovenia, Tiezheng Yu, Willy Chung, and Pascale Fung. 2023. InstructAlign: High-and-low resource language alignment via continual crosslingual instruction tuning. In Proceedings of the First Workshop in South East Asian Language Processing, pages 55–78, Nusa Dua, Bali, Indonesia. Association for Computational Linguistics.
  • Conover et al. (2023) Mike Conover, Matt Hayes, Ankit Mathur, Jianwei Xie, Jun Wan, Sam Shah, Ali Ghodsi, Patrick Wendell, Matei Zaharia, and Reynold Xin. 2023. Free dolly: Introducing the world’s first truly open instruction-tuned llm. Last accessed 2024-01-15.
@@
  • DeepMind (2024) Google DeepMind. 2024. Gemini 1.5: Scaling multimodal ai. https://deepmind.google/technologies/gemini.
@@
  • Gao et al. (2024) Leo Gao, Jonathan Tow, Baber Abbasi, Stella Biderman, Sid Black, Anthony DiPofi, Charles Foster, Laurence Golding, Jeffrey Hsu, Alain Le Noac’h, Haonan Li, Kyle McDonell, Niklas Muennighoff, Chris Ociepa, Jason Phang, Laria Reynolds, Hailey Schoelkopf, Aviya Skowron, Lintang Sutawika, Eric Tang, Anish Thite, Ben Wang, Kevin Wang, and Andy Zou. 2024. A framework for few-shot language model evaluation.
@@
  • Honovich et al. (2023) Or Honovich, Thomas Scialom, Omer Levy, and Timo Schick. 2023. Unnatural instructions: Tuning language models with (almost) no human labor. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 14409–14428, Toronto, Canada. Association for Computational Linguistics.
@@
  • Joshi et al. (2020) Pratik Joshi, Sebastin Santy, Amar Budhiraja, Kalika Bali, and Monojit Choudhury. 2020. The state and fate of linguistic diversity and inclusion in the NLP world. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 6282–6293, Online. Association for Computational Linguistics.
  • Köpf et al. (2023) Andreas Köpf, Yannic Kilcher, Dimitri von Rütte, Sotiris Anagnostidis, Zhi-Rui Tam, Keith Stevens, Abdullah Barhoum, Arnav Dantuluri, et al. 2023. Openassistant conversations – democratizing large language model alignment. arXiv preprint arXiv:2304.07327.
@@
  • Köksal et al. (2024) Abdullatif Köksal, Marion Thaler, Ayyoob Imani, Ahmet Üstün, Anna Korhonen, and Hinrich Schütze. 2024. Muri: High-quality instruction tuning datasets for low-resource languages via reverse instructions.
@@
  • Li et al. (2024) Chong Li, Wen Yang, Jiajun Zhang, Jinliang Lu, Shaonan Wang, and Chengqing Zong. 2024. X-instruction: Aligning language model in low-resource languages with self-curated cross-lingual instructions. In Findings of the Association for Computational Linguistics: ACL 2024, pages 546–566, Bangkok, Thailand. Association for Computational Linguistics.
@@
  • Li et al. (2023) Haonan Li, Fajri Koto, Minghao Wu, Alham Fikri Aji, and Timothy Baldwin. 2023. Bactrian-x: Multilingual replicable instruction-following models with low-rank adaptation. arXiv preprint arXiv:2305.15011.
@@
  • Lin (2004) Chin-Yew Lin. 2004. ROUGE: A package for automatic evaluation of summaries. In Text Summarization Branches Out, pages 74–81, Barcelona, Spain. Association for Computational Linguistics.
@@
  • Liu et al. (2022) Alisa Liu, Swabha Swayamdipta, Noah A. Smith, and Yejin Choi. 2022. WANLI: Worker and AI collaboration for natural language inference dataset creation. In Findings of the Association for Computational Linguistics: EMNLP 2022, pages 6826–6847, Abu Dhabi, United Arab Emirates. Association for Computational Linguistics.
@@
  • Longpre et al. (2023) Shayne Longpre, Le Hou, Tu Vu, Albert Webson, Hyung Won Chung, Yi Tay, Denny Zhou, Quoc V. Le, Barret Zoph, Jason Wei, and Adam Roberts. 2023. The flan collection: Designing data and methods for effective instruction tuning. In Proceedings of the 40th International Conference on Machine Learning, ICML’23. JMLR.org.
@@
  • Mussakhojayeva et al. (2024) Saida Mussakhojayeva, Rinat Gilmullin, Bulat Khakimov, Mansur Galimov, Daniil Orel, Adal Abilbekov, and Huseyin Atakan Varol. 2024. Noise-robust multilingual speech recognition and the tatar speech corpus. In 2024 International Conference on Artificial Intelligence in Information and Communication (ICAIIC), pages 732–737.
@@
  • Mussakhojayeva et al. (2021) Saida Mussakhojayeva, Aigerim Janaliyeva, Almas Mirzakhmetov, Yerbolat Khassanov, and Huseyin Atakan Varol. 2021. KazakhTTS: An Open-Source Kazakh Text-to-Speech Synthesis Dataset. In Proc. Interspeech 2021, pages 2786–2790.
@@
  • Mussakhojayeva et al. (2022) Saida Mussakhojayeva, Yerbolat Khassanov, and Huseyin Atakan Varol. 2022. Ksc2: An industrial-scale open-source kazakh speech corpus. In Interspeech 2022, pages 1367–1371.
@@
  • Nations (n.d.) United Nations. n.d. Un e-government development index (egdi). https://publicadministration.un.org/egovkb/Data-Center. Accessed: 2025-01-19.
  • OpenAI (2024) OpenAI. 2024. Gpt-4o technical report. https://openai.com/research/gpt-4o.
@@
  • Ouyang et al. (2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022. Training language models to follow instructions with human feedback. arXiv preprint arXiv:2203.02155.
  • Qwen et al. (2025) Qwen, :, An Yang, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, Haoran Wei, Huan Lin, Jian Yang, Jianhong Tu, Jianwei Zhang, Jianxin Yang, Jiaxi Yang, Jingren Zhou, Junyang Lin, Kai Dang, Keming Lu, Keqin Bao, Kexin Yang, Le Yu, Mei Li, Mingfeng Xue, Pei Zhang, Qin Zhu, Rui Men, Runji Lin, Tianhao Li, Tianyi Tang, Tingyu Xia, Xingzhang Ren, Xuancheng Ren, Yang Fan, Yang Su, Yichang Zhang, Yu Wan, Yuqiong Liu, Zeyu Cui, Zhenru Zhang, and Zihan Qiu. 2025. Qwen2.5 technical report.
@@
  • Sagyndyk et al. (2024a) Beksultan Sagyndyk, Sanzhar Murzakhmetov, Sanzhar Umbet, and Kirill Yakunin. 2024a. Kazakh constitution: Multiple choice benchmark. Available on Hugging Face.
@@
  • Sagyndyk et al. (2024b) Beksultan Sagyndyk, Sanzhar Murzakhmetov, Sanzhar Umbet, and Kirill Yakunin. 2024b. Kazakh dastur: Multiple choice benchmark. Available on Hugging Face.
@@
  • Sagyndyk et al. (2024c) Beksultan Sagyndyk, Sanzhar Murzakhmetov, Sanzhar Umbet, and Kirill Yakunin. 2024c. Kazakh unified national testing: Multiple choice benchmark. Available on Hugging Face.
@@
  • Sanh et al. (2021) Victor Sanh, Albert Webson, Colin Raffel, Stephen H. Bach, Lintang Sutawika, Zaid Alyafeai, Antoine Chaffin, Arnaud Stiegler, Teven Le Scao, Arun Raja, Manan Dey, M Saiful Bari, Canwen Xu, Urmish Thakker, Shanya Sharma, Eliza Szczechla, Taewoon Kim, Gunjan Chhablani, Nihal Nayak, Debajyoti Datta, Jonathan Chang, Mike Tian-Jian Jiang, Han Wang, Matteo Manica, Sheng Shen, Zheng Xin Yong, Harshit Pandey, Rachel Bawden, Thomas Wang, Trishala Neeraj, Jos Rozen, Abheesht Sharma, Andrea Santilli, Thibault Fevry, Jason Alan Fries, Ryan Teehan, Tali Bers, Stella Biderman, Leo Gao, Thomas Wolf, and Alexander M. Rush. 2021. Multitask prompted training enables zero-shot task generalization. arXiv preprint arXiv:2110.08207.
@@
  • Sengupta et al. (2023) Neha Sengupta, Sunil Kumar Sahu, Bokang Jia, Satheesh Katipomu, Haonan Li, Fajri Koto, Osama Mohammed Afzal, Samta Kamboj, Onkar Pandit, Rahul Pal, Lalit Pradhan, Zain Muhammad Mujahid, Massa Baali, Alham Fikri Aji, Zhengzhong Liu, Andy Hock, Andrew Feldman, Jonathan Lee, Andrew Jackson, Preslav Nakov, Timothy Baldwin, and Eric Xing. 2023. Jais and jais-chat: Arabic-centric foundation and instruction-tuned open generative large language models.
@@
  • Taori et al. (2023) Rohan Taori, Ishaan Gulrajani, Tianyi Zhang, Yann Dubois, Xuechen Li, Carlos Guestrin, Percy Liang, and Tatsunori B. Hashimoto. 2023. Stanford alpaca: An instruction-following llama model. Last accessed 2024-01-15.
@@
  • Team (2024) Falcon-LLM Team. 2024. The falcon 3 family of open models.
@@
  • Team et al. (2024) Gemma Team, Morgane Riviere, Shreya Pathak, Pier Giuseppe Sessa, Cassidy Hardin, Surya Bhupatiraju, Léonard Hussenot, Thomas Mesnard, Bobak Shahriari, Alexandre Ramé, Johan Ferret, Peter Liu, Pouya Tafti, Abe Friesen, Michelle Casbon, Sabela Ramos, Ravin Kumar, Charline Le Lan, Sammy Jerome, Anton Tsitsulin, Nino Vieillard, Piotr Stanczyk, Sertan Girgin, Nikola Momchev, Matt Hoffman, Shantanu Thakoor, Jean-Bastien Grill, Behnam Neyshabur, Olivier Bachem, Alanna Walton, Aliaksei Severyn, Alicia Parrish, Aliya Ahmad, Allen Hutchison, Alvin Abdagic, Amanda Carl, Amy Shen, Andy Brock, Andy Coenen, Anthony Laforge, Antonia Paterson, Ben Bastian, Bilal Piot, Bo Wu, Brandon Royal, Charlie Chen, Chintu Kumar, Chris Perry, Chris Welty, Christopher A. Choquette-Choo, Danila Sinopalnikov, David Weinberger, Dimple Vijaykumar, Dominika Rogozińska, Dustin Herbison, Elisa Bandy, Emma Wang, Eric Noland, Erica Moreira, Evan Senter, Evgenii Eltyshev, Francesco Visin, Gabriel Rasskin, Gary Wei, Glenn Cameron, Gus Martins, Hadi Hashemi, Hanna Klimczak-Plucińska, Harleen Batra, Harsh Dhand, Ivan Nardini, Jacinda Mein, Jack Zhou, James Svensson, Jeff Stanway, Jetha Chan, Jin Peng Zhou, Joana Carrasqueira, Joana Iljazi, Jocelyn Becker, Joe Fernandez, Joost van Amersfoort, Josh Gordon, Josh Lipschultz, Josh Newlan, Ju yeong Ji, Kareem Mohamed, Kartikeya Badola, Kat Black, Katie Millican, Keelin McDonell, Kelvin Nguyen, Kiranbir Sodhia, Kish Greene, Lars Lowe Sjoesund, Lauren Usui, Laurent Sifre, Lena Heuermann, Leticia Lago, Lilly McNealus, Livio Baldini Soares, Logan Kilpatrick, Lucas Dixon, Luciano Martins, Machel Reid, Manvinder Singh, Mark Iverson, Martin Görner, Mat Velloso, Mateo Wirth, Matt Davidow, Matt Miller, Matthew Rahtz, Matthew Watson, Meg Risdal, Mehran Kazemi, Michael Moynihan, Ming Zhang, Minsuk Kahng, Minwoo Park, Mofi Rahman, Mohit Khatwani, Natalie Dao, Nenshad Bardoliwalla, Nesh Devanathan, Neta Dumai, Nilay Chauhan, Oscar Wahltinez, Pankil Botarda, Parker Barnes, Paul Barham, Paul Michel, Pengchong Jin, Petko Georgiev, Phil Culliton, Pradeep Kuppala, Ramona Comanescu, Ramona Merhej, Reena Jana, Reza Ardeshir Rokni, Rishabh Agarwal, Ryan Mullins, Samaneh Saadat, Sara Mc Carthy, Sarah Cogan, Sarah Perrin, Sébastien M. R. Arnold, Sebastian Krause, Shengyang Dai, Shruti Garg, Shruti Sheth, Sue Ronstrom, Susan Chan, Timothy Jordan, Ting Yu, Tom Eccles, Tom Hennigan, Tomas Kocisky, Tulsee Doshi, Vihan Jain, Vikas Yadav, Vilobh Meshram, Vishal Dharmadhikari, Warren Barkley, Wei Wei, Wenming Ye, Woohyun Han, Woosuk Kwon, Xiang Xu, Zhe Shen, Zhitao Gong, Zichuan Wei, Victor Cotruta, Phoebe Kirk, Anand Rao, Minh Giang, Ludovic Peran, Tris Warkentin, Eli Collins, Joelle Barral, Zoubin Ghahramani, Raia Hadsell, D. Sculley, Jeanine Banks, Anca Dragan, Slav Petrov, Oriol Vinyals, Jeff Dean, Demis Hassabis, Koray Kavukcuoglu, Clement Farabet, Elena Buchatskaya, Sebastian Borgeaud, Noah Fiedel, Armand Joulin, Kathleen Kenealy, Robert Dadashi, and Alek Andreev. 2024. Gemma 2: Improving open language models at a practical size.
@@
  • Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971.
@@
  • Wang et al. (2024) Renxi Wang, Haonan Li, Minghao Wu, Yuxia Wang, Xudong Han, Chiyu Zhang, and Timothy Baldwin. 2024. Demystifying instruction mixing for fine-tuning large language models. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 4: Student Research Workshop), pages 68–75, Bangkok, Thailand. Association for Computational Linguistics.
@@
  • Wang et al. (2023) Yizhong Wang, Yeganeh Kordi, Swaroop Mishra, Alisa Liu, Noah A. Smith, Daniel Khashabi, and Hannaneh Hajishirzi. 2023. Self-instruct: Aligning language models with self-generated instructions. arXiv preprint arXiv:2212.10560.

### v3

@@ 6 Conclusion and Future Work
The evaluation results show that this approach substantially improved the model’s factual knowledge and the understanding of low-resource languages. It also shows that after such fine-tuning, the model’s responses are much better in terms of correctness and soundness, as assessed by native speakers and LLM as a judge.
In future work, we plan to apply this methodology to other languages and dialects. We further aim to work towards streamlining and automating the process as much as possible. We will also focus more on the modeling part of the experiments, and open-source culturally and institutionally relevant models for low-resource languages, including Kazakh.
@@ 8 Ethics
  • Z. Alyafeai, K. Almubarak, A. Ashraf, D. Alnuhait, S. Alshahrani, G. Abdulrahman, G. Ahmed, Q. Gawah, Z. Saleh, M. Ghaleb, Y. Ali, and M. Al-shaibani (2024) CIDAR: culturally relevant instruction dataset for arabic. In Findings of the Association for Computational Linguistics: ACL 2024, Bangkok, Thailand, pp. 12878–12901. External Links: Link, Document Cited by: §2.2.
  • Anthropic (2024) Claude ai: an overview. Note: https://www.anthropic.com Cited by: §4.2.
  • L. Bandarkar, D. Liang, B. Muller, M. Artetxe, S. N. Shukla, D. Husa, N. Goyal, A. Krishnan, L. Zettlemoyer, and M. Khabsa (2024) The belebele benchmark: a parallel reading comprehension dataset in 122 language variants. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), L. Ku, A. Martins, and V. Srikumar (Eds.), Bangkok, Thailand, pp. 749–775. External Links: Link, Document Cited by: §2.3.
@@
  • M. Brief, O. Ovadia, G. Shenderovitz, N. B. Yoash, R. Lemberg, and E. Sheetrit (2024) Mixing it up: the cocktail effect of multi-task fine-tuning on llm performance – a case study in finance. External Links: 2410.01109, Link Cited by: §5.1.
@@
  • S. Cahyawijaya, H. Lovenia, F. Koto, R. Putri, W. Cenggoro, J. Lee, S. Akbar, E. Dave, N. Nuurshadieq, M. Mahendra, R. Putri, B. Wilie, G. Winata, A. Aji, A. Purwarianti, and P. Fung (2024) Cendol: open instruction-tuned generative large language models for Indonesian languages. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), L. Ku, A. Martins, and V. Srikumar (Eds.), Bangkok, Thailand, pp. 14899–14914. External Links: Link, Document Cited by: §1.
@@
  • S. Cahyawijaya, H. Lovenia, T. Yu, W. Chung, and P. Fung (2023) InstructAlign: high-and-low resource language alignment via continual crosslingual instruction tuning. In Proceedings of the First Workshop in South East Asian Language Processing, D. Wijaya, A. F. Aji, C. Vania, G. I. Winata, and A. Purwarianti (Eds.), Nusa Dua, Bali, Indonesia, pp. 55–78. External Links: Link, Document Cited by: §1, §2.2.
  • M. Conover, M. Hayes, A. Mathur, J. Xie, J. Wan, S. Shah, A. Ghodsi, P. Wendell, M. Zaharia, and R. Xin (2023) External Links: Link Cited by: §1, §1, §2.1.
@@
  • G. DeepMind (2024) Gemini 1.5: scaling multimodal ai. Note: https://deepmind.google/technologies/gemini Cited by: §4.2.
@@
  • L. Gao, J. Tow, B. Abbasi, S. Biderman, S. Black, A. DiPofi, C. Foster, L. Golding, J. Hsu, A. Le Noac’h, H. Li, K. McDonell, N. Muennighoff, C. Ociepa, J. Phang, L. Reynolds, H. Schoelkopf, A. Skowron, L. Sutawika, E. Tang, A. Thite, B. Wang, K. Wang, and A. Zou (2024) A framework for few-shot language model evaluation. Zenodo. External Links: Document, Link Cited by: §5.1.
@@
  • O. Honovich, T. Scialom, O. Levy, and T. Schick (2023) Unnatural instructions: tuning language models with (almost) no human labor. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), A. Rogers, J. Boyd-Graber, and N. Okazaki (Eds.), Toronto, Canada, pp. 14409–14428. External Links: Link, Document Cited by: §2.1.
@@
  • P. Joshi, S. Santy, A. Budhiraja, K. Bali, and M. Choudhury (2020) The state and fate of linguistic diversity and inclusion in the NLP world. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, D. Jurafsky, J. Chai, N. Schluter, and J. Tetreault (Eds.), Online, pp. 6282–6293. External Links: Link, Document Cited by: §1.
@@
  • A. Köksal, M. Thaler, A. Imani, A. Üstün, A. Korhonen, and H. Schütze (2024) MURI: high-quality instruction tuning datasets for low-resource languages via reverse instructions. External Links: 2409.12958 Cited by: §2.2, §4.2.
  • A. Köpf, Y. Kilcher, D. Von Rütte, S. Anagnostidis, Z. R. Tam, K. Stevens, A. Barhoum, D. Nguyen, O. Stanley, R. Nagyfi, et al. (2023) Openassistant conversations-democratizing large language model alignment. Advances in Neural Information Processing Systems 36, pp. 47669–47681. Cited by: §2.1.
@@
  • C. Li, W. Yang, J. Zhang, J. Lu, S. Wang, and C. Zong (2024) X-instruction: aligning language model in low-resource languages with self-curated cross-lingual instructions. In Findings of the Association for Computational Linguistics: ACL 2024, L. Ku, A. Martins, and V. Srikumar (Eds.), Bangkok, Thailand, pp. 546–566. External Links: Link, Document Cited by: §2.2.
@@
  • H. Li, F. Koto, M. Wu, A. F. Aji, and T. Baldwin (2023) Bactrian-x: multilingual replicable instruction-following models with low-rank adaptation. arXiv preprint arXiv:2305.15011. Cited by: §1, §1, §2.2.
@@
  • C. Lin (2004) ROUGE: a package for automatic evaluation of summaries. In Text Summarization Branches Out, Barcelona, Spain, pp. 74–81. External Links: Link Cited by: §5.2.
@@
  • A. Liu, S. Swayamdipta, N. A. Smith, and Y. Choi (2022) WANLI: worker and AI collaboration for natural language inference dataset creation. In Findings of the Association for Computational Linguistics: EMNLP 2022, Y. Goldberg, Z. Kozareva, and Y. Zhang (Eds.), Abu Dhabi, United Arab Emirates, pp. 6826–6847. External Links: Link, Document Cited by: §1.
@@
  • S. Longpre, L. Hou, T. Vu, A. Webson, H. W. Chung, Y. Tay, D. Zhou, Q. V. Le, B. Zoph, J. Wei, and A. Roberts (2023) The flan collection: designing data and methods for effective instruction tuning. In Proceedings of the 40th International Conference on Machine Learning, ICML’23, Cited by: §1, §2.1.
@@
  • S. Mussakhojayeva, R. Gilmullin, B. Khakimov, M. Galimov, D. Orel, A. Abilbekov, and H. A. Varol (2024) Noise-robust multilingual speech recognition and the tatar speech corpus. In 2024 International Conference on Artificial Intelligence in Information and Communication (ICAIIC), Vol. , pp. 732–737. External Links: Document Cited by: §2.3.
@@
  • S. Mussakhojayeva, A. Janaliyeva, A. Mirzakhmetov, Y. Khassanov, and H. A. Varol (2021) KazakhTTS: An Open-Source Kazakh Text-to-Speech Synthesis Dataset. In Proc. Interspeech 2021, pp. 2786–2790. External Links: Document Cited by: §2.3.
@@
  • S. Mussakhojayeva, Y. Khassanov, and H. Atakan Varol (2022) KSC2: an industrial-scale open-source kazakh speech corpus. In Interspeech 2022, pp. 1367–1371. External Links: Document, ISSN 2958-1796 Cited by: §2.3.
@@
  • U. Nations (n.d.) UN e-government development index (egdi). Note: https://publicadministration.un.org/egovkb/Data-CenterAccessed: 2025-01-19 Cited by: §3.
  • OpenAI (2024) GPT-4o technical report. Note: https://openai.com/research/gpt-4o Cited by: §1, §4.2.
@@
  • L. Ouyang, J. Wu, X. Jiang, D. Almeida, C. L. Wainwright, P. Mishkin, C. Zhang, S. Agarwal, K. Slama, A. Ray, et al. (2022) Training language models to follow instructions with human feedback. In Advances in Neural Information Processing Systems, Cited by: §1.
  • Qwen, :, A. Yang, B. Yang, B. Zhang, B. Hui, B. Zheng, B. Yu, C. Li, D. Liu, F. Huang, H. Wei, H. Lin, J. Yang, J. Tu, J. Zhang, J. Yang, J. Yang, J. Zhou, J. Lin, K. Dang, K. Lu, K. Bao, K. Yang, L. Yu, M. Li, M. Xue, P. Zhang, Q. Zhu, R. Men, R. Lin, T. Li, T. Tang, T. Xia, X. Ren, X. Ren, Y. Fan, Y. Su, Y. Zhang, Y. Wan, Y. Liu, Z. Cui, Z. Zhang, and Z. Qiu (2025) Qwen2.5 technical report. External Links: 2412.15115, Link Cited by: §5.
@@
  • B. Sagyndyk, S. Murzakhmetov, S. Umbet, and K. Yakunin (2024a) Kazakh constitution: multiple choice benchmark. Note: Available on Hugging Face External Links: Link Cited by: §5.1.
@@
  • B. Sagyndyk, S. Murzakhmetov, S. Umbet, and K. Yakunin (2024b) Kazakh dastur: multiple choice benchmark. Note: Available on Hugging Face External Links: Link Cited by: §5.1.
@@
  • B. Sagyndyk, S. Murzakhmetov, S. Umbet, and K. Yakunin (2024c) Kazakh unified national testing: multiple choice benchmark. Note: Available on Hugging Face External Links: Link Cited by: §5.1.
  • B. Sagyndyk, S. Murzakhmetov, and K. Yakunin (2025) Kaz-RoBERTa conversational technical report. TechRxiv. External Links: Document, Link Cited by: §5.2.
@@
  • V. Sanh, A. Webson, C. Raffel, S. H. Bach, L. Sutawika, Z. Alyafeai, A. Chaffin, A. Stiegler, T. L. Scao, A. Raja, M. Dey, M. S. Bari, C. Xu, U. Thakker, S. Sharma, E. Szczechla, T. Kim, G. Chhablani, N. Nayak, D. Datta, J. Chang, M. T. Jiang, H. Wang, M. Manica, S. Shen, Z. X. Yong, H. Pandey, R. Bawden, T. Wang, T. Neeraj, J. Rozen, A. Sharma, A. Santilli, T. Fevry, J. A. Fries, R. Teehan, T. Bers, S. Biderman, L. Gao, T. Wolf, and A. M. Rush (2021) Multitask prompted training enables zero-shot task generalization. In International Conference on Learning Representations, Cited by: §1, §2.1.
@@
  • N. Sengupta, S. K. Sahu, B. Jia, S. Katipomu, H. Li, F. Koto, O. M. Afzal, S. Kamboj, O. Pandit, R. Pal, L. Pradhan, Z. M. Mujahid, M. Baali, A. F. Aji, Z. Liu, A. Hock, A. Feldman, J. Lee, A. Jackson, P. Nakov, T. Baldwin, and E. Xing (2023) Jais and jais-chat: arabic-centric foundation and instruction-tuned open generative large language models. External Links: 2308.16149 Cited by: §1, §2.2.
@@
  • R. Taori, I. Gulrajani, T. Zhang, Y. Dubois, X. Li, C. Guestrin, P. Liang, and T. B. Hashimoto (2023) Stanford alpaca: an instruction-following llama model. Note: Last accessed 2024-01-15 Cited by: §2.2.
@@
  • F. Team (2024) The falcon 3 family of open models. External Links: Link Cited by: §5.
@@
  • G. Team, M. Riviere, S. Pathak, P. G. Sessa, C. Hardin, S. Bhupatiraju, L. Hussenot, T. Mesnard, B. Shahriari, A. Ramé, J. Ferret, P. Liu, P. Tafti, A. Friesen, M. Casbon, S. Ramos, R. Kumar, C. L. Lan, S. Jerome, A. Tsitsulin, N. Vieillard, P. Stanczyk, S. Girgin, N. Momchev, M. Hoffman, S. Thakoor, J. Grill, B. Neyshabur, O. Bachem, A. Walton, A. Severyn, A. Parrish, A. Ahmad, A. Hutchison, A. Abdagic, A. Carl, A. Shen, A. Brock, A. Coenen, A. Laforge, A. Paterson, B. Bastian, B. Piot, B. Wu, B. Royal, C. Chen, C. Kumar, C. Perry, C. Welty, C. A. Choquette-Choo, D. Sinopalnikov, D. Weinberger, D. Vijaykumar, D. Rogozińska, D. Herbison, E. Bandy, E. Wang, E. Noland, E. Moreira, E. Senter, E. Eltyshev, F. Visin, G. Rasskin, G. Wei, G. Cameron, G. Martins, H. Hashemi, H. Klimczak-Plucińska, H. Batra, H. Dhand, I. Nardini, J. Mein, J. Zhou, J. Svensson, J. Stanway, J. Chan, J. P. Zhou, J. Carrasqueira, J. Iljazi, J. Becker, J. Fernandez, J. van Amersfoort, J. Gordon, J. Lipschultz, J. Newlan, J. Ji, K. Mohamed, K. Badola, K. Black, K. Millican, K. McDonell, K. Nguyen, K. Sodhia, K. Greene, L. L. Sjoesund, L. Usui, L. Sifre, L. Heuermann, L. Lago, L. McNealus, L. B. Soares, L. Kilpatrick, L. Dixon, L. Martins, M. Reid, M. Singh, M. Iverson, M. Görner, M. Velloso, M. Wirth, M. Davidow, M. Miller, M. Rahtz, M. Watson, M. Risdal, M. Kazemi, M. Moynihan, M. Zhang, M. Kahng, M. Park, M. Rahman, M. Khatwani, N. Dao, N. Bardoliwalla, N. Devanathan, N. Dumai, N. Chauhan, O. Wahltinez, P. Botarda, P. Barnes, P. Barham, P. Michel, P. Jin, P. Georgiev, P. Culliton, P. Kuppala, R. Comanescu, R. Merhej, R. Jana, R. A. Rokni, R. Agarwal, R. Mullins, S. Saadat, S. M. Carthy, S. Cogan, S. Perrin, S. M. R. Arnold, S. Krause, S. Dai, S. Garg, S. Sheth, S. Ronstrom, S. Chan, T. Jordan, T. Yu, T. Eccles, T. Hennigan, T. Kocisky, T. Doshi, V. Jain, V. Yadav, V. Meshram, V. Dharmadhikari, W. Barkley, W. Wei, W. Ye, W. Han, W. Kwon, X. Xu, Z. Shen, Z. Gong, Z. Wei, V. Cotruta, P. Kirk, A. Rao, M. Giang, L. Peran, T. Warkentin, E. Collins, J. Barral, Z. Ghahramani, R. Hadsell, D. Sculley, J. Banks, A. Dragan, S. Petrov, O. Vinyals, J. Dean, D. Hassabis, K. Kavukcuoglu, C. Farabet, E. Buchatskaya, S. Borgeaud, N. Fiedel, A. Joulin, K. Kenealy, R. Dadashi, and A. Andreev (2024) Gemma 2: improving open language models at a practical size. External Links: 2408.00118, Link Cited by: §5.
@@
  • H. Touvron, T. Lavril, G. Izacard, X. Martinet, M. Lachaux, T. Lacroix, B. Rozière, N. Goyal, E. Hambro, F. Azhar, et al. (2023) LLaMA: open and efficient foundation language models. arXiv preprint arXiv:2302.13971. Cited by: §4.2.
@@
  • R. Wang, H. Li, M. Wu, Y. Wang, X. Han, C. Zhang, and T. Baldwin (2024) Demystifying instruction mixing for fine-tuning large language models. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 4: Student Research Workshop), X. Fu and E. Fleisig (Eds.), Bangkok, Thailand, pp. 68–75. External Links: Link, Document Cited by: §5.1.
@@
  • Y. Wang, Y. Kordi, S. Mishra, A. Liu, N. A. Smith, D. Khashabi, and H. Hajishirzi (2023) Self-instruct: aligning language models with self-generated instructions. arXiv preprint arXiv:2212.10560. External Links: Link Cited by: §2.1.